### v2

# Probing metric fluctuations with the spin decoherence of a particle in a quantum simulation

Abstract
Exploring potential empirical manifestations of quantum gravity is a challenging pursuit. In this study, we utilise a lattice representation of a (2+1)(2+1)D massive gravity toy model interacting with Dirac fermions that can support specific spacetime fluctuations. We focus on the evolution of the fermion’s spin due to its coupling to spacetime fluctuations. To monitor their dynamics a minimal model is required that comprises two bosonic modes describing spacetime geometry fluctuations coupled with the spin of the fermion. A possible emulation of this system involves encoding spin degrees of freedom in the electronic states of an atom coupled with a bimodal optical cavity that provides the two bosonic modes. The spin exhibits a variety of dynamical behaviours due to its coupling with the fluctuating geometry, with decoherence emerging as a key signature of this interaction. Our proposal introduces a novel approach for modelling the effect of interactions between quantum gravity and matter that can be probed with current technology.
articletype:
1 Introduction
Understanding the fundamental nature of spacetime at quantum scales is a central challenge in contemporary theoretical physics. Quantum gravity, the sought-after theory unifying general relativity and quantum physics, remains an elusive theory, with several theoretical and empirical uncertainties. Apart from its conceptual difficulties this theory is strongly interacting and thus often too complex to analytically or numerically investigate. Despite significant advances in theoretical frameworks like lower-dimensional quantum gravity, string theory and loop quantum gravity, observations and empirical validations of quantum gravitational effects remains well beyond the reach of our current technologies [1, 2].
An interesting possibility to probe gravitational degrees of freedom and study properties that are otherwise inaccessible is through the lens of condensed matter physics, where collective excitations that resemble gravitons have been recently found. For example, massive gravitons have been shown to emerge in the 3He-B superfluid, where gravitational effects arise from fermionic bilinears following symmetry breaking [3]. This bears a compelling resemblance to Diakonov’s proposal for a lattice-regularized quantum gravity theory, which involves a massless emergent gravitational field [4].
Another significant area of exploration involves certain collective excitations in fractional quantum Hall liquids, which remarkably behave as analogue gravitons. Specifically, the Girvin-MacDonald-Platzman mode of the fractional quantum Hall effect [5, 6], often termed the magnetoroton, is recognized as a massive graviton excitation in its long-wavelength limit [7, 8, 9]. Moreover, the predicted properties of these excitations are in agreement with current experimental observations [10, 11, 12]. However, the strong interactions inherent to fractional quantum Hall liquids make them analytically intractable, highlighting the need for alternative approaches that can emulate their emergent features.
To overcome these analytical limitations, an alternative strategy involves engineering systems capable of effectively simulating quantum signatures of gravity. Such quantum simulations offer the advantage of accessing a broad spectrum of coupling regimes. In particular, they enable to amplify properties of gravity that are otherwise experimentally inaccessible due to its very weak coupling to matter. Recent progress in simulating quantum gauge theories, for instance, has provided valuable insights into complex phenomena like quark confinement [13]. While various platforms exist for simulating classical gravity, such as deforming graphene-like systems to produce non-trivial extrinsic geometry [14, 15] or creating tunneling coupling inhomogeneities that generate intrinsic geometries [16, 17], simulating quantum gravity has remained an open challenge. Roadblocks persist at both conceptual levels, like realizing quantum fluctuations of spacetime, and practical levels, such as engineering the intricate self-interactions characteristic of gravitational theories.
This work is motivated by the importance of bridging the gap between theory and experiment in quantum gravity. We propose an experimental setup that can simulate quantum metric fluctuations, thus facilitating the identification of novel physical manifestations of quantum gravity. Our starting point is based on a lattice representation of a particular model for metric perturbations in $(2+1)$ dimensions coupled to Dirac fermions [18]. Unlike the model considered in [18], the fluctuations of the spatial geometry considered here are described by a traceless symmetric tensor, which can be associated to a unimodular spatial metric. Therefore, our lattice model gives rise to non-relativistic massive metric perturbations similar to the massive gravitons identified in the study of the Girvin–MacDonald–Platzman (GMP) mode in the fractional quantum Hall effect [8, 9]. This framework allows us to focus on a simplified set of spacetime fluctuations that still exhibit a rich dynamical behaviour. We employ the interactions between matter and spacetime fluctuations to use the spin of the Dirac fermions as a probe to deduce the behaviour of the gravitational dynamics.
The lattice representation of quantum gravity coupled to matter has also the advantage of being readily amenable to experimental simulations, thus bridging the gap between theory and experiment in this field of research. In particular, we identify the smallest instance of the lattice model that encodes the coupling between spacetime fluctuations and the spin of the Dirac fermion. We then propose an experimental setup for the quantum simulation of this minimal model. This setup involves encoding the spin degrees of freedom in the electronic levels of a single atom, placed inside a bimodal optical cavity that encodes two bosonic modes responsible for the spacetime fluctuations. By varying the coupling between the atom and the cavity we are able to use the atom as a probe to identify the dynamics of the simulated spacetime fluctuations. We identify distinct dynamical signatures of the spin across different gravitational regimes, illustrating how spacetime fluctuations influence matter evolution at the quantum level. These findings open a pathway for experimental probes of quantum gravitational phenomena possibly by enhancing the effect fluctuating gravity can have on matter. Moreover, our research establishes a platform for simulating more intricate models of gravity–matter interactions in future work. Beyond their fundamental significance, such quantum simulations could shed light on information flow, entanglement dynamics, and decoherence in curved quantum geometries, offering valuable insights for both quantum gravity and quantum technologies.
2 Lattice representation of Dirac field coupled to a massive spin-2 field
We now consider a massless Dirac fermion field coupled to metric perturbations described by a toy model Hamiltonian for a bosonic massive spin-2 field. Subsequently, we will consider a lattice model that describes this composite system in its low energy limit.
2.1 Carrollian massive spin-2 field
For simplicity we consider purely spatial metric fluctuations $h_{ij}$ around a flat-space background. We adopt Cartesian coordinates $x^{i}=\{x,y\}$ so that the spatial metric tensor has the form
$g_{ij}=\delta_{ij}+h_{ij}.$ (1)
Moreover, we restrict $h_{ij}$ to be traceless i.e., $h\equiv\delta^{ij}h_{ij}=0$. This has the consequence that the metric determinant is fixed to be ${\rm det}(g)=1$, a characteristic of unimodular gravity models. Since Einstein gravity in 2+1 dimensions lacks propagating degrees of freedom [19], the massless Fierz-Pauli action coming from linearising the Einstein-Hilbert Lagrangian has trivial dynamics for $h_{ij}$. To introduce non-trivial dynamics, we instead model $h_{ij}$ as a massive spin-2 field. The adopted Fierz–Pauli action for a relativistic massive spin-2 field in 2+1 dimensions, describes two local propagating degrees of freedom [20], thus having a very similar behaviour to the linearised Einstein gravity in four dimensions. For our purposes, it is sufficient to focus on the temporal dynamics of the metric fluctuations, neglecting spatial derivatives of $h_{ij}$, while still capturing the essential physical behaviour and complexity of the system. We therefore consider the following action principle
$S_{\rm gr}=\frac{1}{64\pi G}\int dt\int d^{2}x\;\bigg(-\dot{h}_{ij}\dot{h}^{ij}-\dot{h}^{2}+\mu^{2}\left(h_{ij}h^{ij}-h^{2}\right)+\xi\,h\bigg),$ (2)
where $\mu$ is a mass parameter and $\xi$ is a Lagrange multiplier enforcing the tracelessness condition. This non-relativistic Hamiltonian for metric perturbations can be obtained as the electric-type Carrollian limit of the Fierz–Pauli Lagrangian for a spin-2 field [21] after integrating out the field components $h_{0i}$ and by identifying $h_{00}=\xi$. By rescaling the metric fluctuations as
$h_{ij}\rightarrow\sqrt{16\pi G}\,h_{ij},$ (3)
the canonical momenta of the model are given by
$p_{ij}=\frac{1}{2}\left(\dot{h}_{ij}-\dot{h}\delta_{ij}\right),\qquad p_{\xi}=0,$ (4)
satysfying canonical Poisson brackets
$\{h_{ij}(\boldsymbol{x}),p^{kl}(\boldsymbol{x}^{\prime})\}=\left(\delta_{i}^{k}\delta_{j}^{l}+\delta_{i}^{l}\delta_{j}^{k}\right)\delta^{(2)}(\boldsymbol{x}-\boldsymbol{x}^{\prime}),\qquad\{\xi(\boldsymbol{x}),p_{\xi}(\boldsymbol{x}^{\prime})\}=\delta^{(2)}(\boldsymbol{x}-\boldsymbol{x}^{\prime}).$ (5)
Dirac’s algorithm for constrained systems leads to second-class constraints given by $\phi_{1}=p_{\xi}$, $\phi_{2}=h$, $\phi_{3}=p$, $\phi_{4}=\xi$, where $p\equiv\delta^{ij}p_{ij}$, which can be eliminated by replacing the Poisson bracket by the Dirac bracket. The resulting reduced phase space action leads to the Hamiltonian
$H_{\rm gr}=\int d^{2}x\;\left(p_{ij}p^{ij}+\frac{\mu^{2}}{4}h_{ij}h^{ij}\right),$ (6)
together with the Dirac bracket
$\{h_{ij}(\boldsymbol{x}),p^{kl}(\boldsymbol{x}^{\prime})\}^{*}=\left(\delta_{i}^{k}\delta_{j}^{l}+\delta_{i}^{l}\delta_{j}^{k}-\delta_{ij}\delta^{kl}\right)\delta^{(2)}(\boldsymbol{x}-\boldsymbol{x}^{\prime}).$ (7)
To quantise the model, $h_{ij}$ and $p_{ij}$ are promoted to operators with commutation relations compatible with the Dirac bracket $\{\,,\,\}^{*}\rightarrow-i[\,$,\,]. Equivalently, we parametrise the independent components of $h_{ij}$ and $p_{ij}$ in terms of operators $\tilde{\alpha}$ and $\tilde{\beta}$ satisfying the commutation relations
$\tilde{\alpha}(\boldsymbol{x}),\tilde{\alpha}^{\dagger}(\boldsymbol{x}^{\prime})]=\delta^{(2)}(\boldsymbol{x}-\boldsymbol{x}^{\prime})=[\tilde{\beta}(\boldsymbol{x}),\tilde{\beta}^{\dagger}(\boldsymbol{x}^{\prime})],$ (8)
as
$h_{11}(\boldsymbol{x})=\frac{1}{\sqrt{2}}\left(\tilde{\alpha}(\boldsymbol{x})+\tilde{\alpha}^{\dagger}(\boldsymbol{x})\right),\qquad h_{12}(\boldsymbol{x})=\frac{1}{\sqrt{2}}\left(\tilde{\beta}(\boldsymbol{x})+\tilde{\beta}^{\dagger}(\boldsymbol{x})\right),$ (9)
$p_{11}(\boldsymbol{x})=\frac{i}{\sqrt{2}}\left(\tilde{\alpha}(\boldsymbol{x})-\tilde{\alpha}^{\dagger}(\boldsymbol{x})\right),\qquad p_{12}(\boldsymbol{x})=\frac{i}{\sqrt{2}}\left(\tilde{\beta}(\boldsymbol{x})-\tilde{\beta}^{\dagger}(\boldsymbol{x})\right).$
In momentum space the Hamiltonian can be written as $H_{\rm gr}=H[\tilde{\alpha}]+H[\tilde{\beta}$] with
$H[\tilde{\alpha}]=2\int d^{2}k\bigg[\left(\frac{\mu^{2}}{4}-1\right)\tilde{\alpha}(\boldsymbol{k})\tilde{\alpha}(-\boldsymbol{k})+\left(\frac{\mu^{2}}{4}+1\right)\tilde{\alpha}(\boldsymbol{k})\tilde{\alpha}^{\dagger}(\boldsymbol{k})+{\rm h.c.}\bigg],$ (10)
and similarly for $H[\tilde{\beta}$]. By applying the Bogoliubov transformation
[…]
with
[…]
the Hamiltonian takes the form
$H_{\rm gr}$ $=2\mu\,\int d^{2}k\bigg[\alpha^{\dagger}(\boldsymbol{k})\alpha(\boldsymbol{k})+\beta^{\dagger}(\boldsymbol{k})\beta(\boldsymbol{k})+\frac{1}{2}\delta(0)\bigg],$ (13)
where the last constant term can be dropped.
2.2 Coupling to Dirac fermion
To describe the coupling of these fluctuations with a Dirac field, we introduce the spatial zweibein $e^{a}_{i}$, with $a=1,2$, such that $g_{ij}=\delta_{ab}e^{a}_{i}e^{b}_{j}$. A possible choice leading to the spatial metric (1) after taking into account the rescaling (3) is
$e^{a}_{i}=\delta^{a}_{i}+2\sqrt{\pi G}\delta^{aj}h_{ij}.$ (14)
The Dirac Hamiltonian on this geometry is given by
$H_{\rm Dirac}=\frac{i}{2}\left(\bar{\psi}e^{i}_{a}\gamma^{a}\partial_{i}\psi-\partial_{i}\bar{\psi}e^{i}_{a}\gamma^{a}\psi\right),$ (15)
where $\gamma^{a}$ are gamma matrices and $e^{i}_{a}$ is the inverse zweibein, satisfying $e^{i}_{a}e^{a}_{j}=\delta^{i}_{j}$, and $e^{i}_{a}e^{b}_{i}=\delta^{b}_{a}$.
The Hamiltonian of the full model is then given by
$H=\int d^{2}x\;\psi^{\dagger}\;h\;\psi+H_{\text{gr}},$ (16)
where the single particle Dirac Hamiltonian $h$ reads
$h=\gamma^{0}\left(\delta^{i}_{a}\gamma^{a}-2\sqrt{\pi G}\delta_{aj}h^{ij}\right)(-i\partial_{i})+\frac{i\sqrt{\pi G}}{2}\delta_{aj}\partial_{i}h^{ij}\gamma^{0}\gamma^{a},$ (17)
describing (2+1)(2+1)D Dirac fermions coupled to semiclassical gravity fluctuations. In the following we will present a lattice model that gives rise to this Hamiltonian in the continuum limit.
2.3 Lattice Hamiltonian
Figure 1: (Left) The brick wall lattice representation of the (2+1)(2+1)D Dirac fermions coupled to spacetime fluctuations given by Hamiltonian (18). The unit cell includes two fermionic modes, $a$ and $b$, that encode the spin of the Dirac field. The bosonic modes $\tilde{\alpha}$ and $\tilde{\beta}$ describe spacetime fluctuations and are coupled to the fermions through their tunnelling. The basis vectors $\boldsymbol{n}_{1}$ and $\boldsymbol{n}_{2}$ span the brick wall lattice of fermions and $J_{x}$, $J_{y}$ and $J_{z}$ are the fermionic tunnelling couplings that depend on the bosonic modes $\tilde{\alpha}$ and $\tilde{\beta}$. (Right) The minimal system considered here with two fermionic modes that encode a single spin and two bosonic modes coupled to the fermions.
The (2+1)(2+1)D Dirac fermion coupled to the massive spin-2 field can effectively arise from a 2D lattice model. In [18] a two-dimensional lattice of interacting bosonic and fermionic modes is presented that gives rise to a Hamiltonian that resembles (16) in its low-energy limit. A similar strategy is adopted here where we start with a brick wall lattice of fermions, as shown in Fig. 1 (Left). The lattice Hamiltonian $H_{\rm latt}$ is given by
$H_{\rm latt}=H_{\text{f}+\text{b}}+H_{\text{b}},$ (18)
where the fermionic and bosonic term $H_{\text{f}+\text{b}}$ gives rise to Dirac fermions coupled to the gravity fluctuations, while the bosonic term, $H_{\text{b}}$, corresponds to the pure gravity term. The gravitational lattice Hamiltonian is given by
$H_{\rm b}=H[\tilde{\alpha}]+H[\tilde{\beta}$], (19)
where
$H[\tilde{\alpha}]=\sum_{\boldsymbol{i}}\bigg(\frac{\mu^{2}}{2}(\tilde{\alpha}_{\boldsymbol{i}}+\tilde{\alpha}^{\dagger}_{\boldsymbol{i}})(\tilde{\alpha}_{\boldsymbol{i}}+\tilde{\alpha}^{\dagger}_{\boldsymbol{i}})-2(\tilde{\alpha}_{\boldsymbol{i}}-\tilde{\alpha}^{\dagger}_{\boldsymbol{i}})(\tilde{\alpha}_{\boldsymbol{i}}-\tilde{\alpha}^{\dagger}_{\boldsymbol{i}})\bigg).$ (20)
The fermion-boson interaction term in (18) describes a two-dimensional brick wall lattice, where fermionic operators, $a$ and $b$, live at its vertices and the bosons couple with them when the fermions tunnel to neighboring sites [18]. These fermions satisfy
$a^{\dagger}_{\boldsymbol{i}},a_{\boldsymbol{j}}]=2\delta_{\boldsymbol{i},\boldsymbol{j}}=[b^{\dagger}_{\boldsymbol{i}},b_{\boldsymbol{j}}],$ (21)
and are subject to the tunnelling Hamiltonian
$H_{\text{f-b}}=\sum_{\boldsymbol{i}}\bigg(a^{\dagger}_{\boldsymbol{i}}J_{X}(\boldsymbol{i})b_{\boldsymbol{i}+\boldsymbol{n}_{1}}+a^{\dagger}_{\boldsymbol{i}}J_{Y}(\boldsymbol{i})b_{\boldsymbol{i}+\boldsymbol{n}_{2}}+a^{\dagger}_{\boldsymbol{i}}J_{Z}(\boldsymbol{i})b_{\boldsymbol{i}}\bigg)+{\rm h.c.},$ (22)
where $\boldsymbol{i}=(i_{x},i_{y})$ gives the position of unit cells on the lattice and the vectors
$\boldsymbol{n}_{1}=\left(-\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}}\right),\qquad\boldsymbol{n}_{2}=\left(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}}\right),$ (23)
transport between unit cells. The tunnelling couplings $J_{\alpha}(\boldsymbol{i})$, $\alpha=X,Y,Z$ depend on the bosonic fluctuations $\tilde{\alpha}_{\boldsymbol{i}}$ and $\tilde{\beta}_{\boldsymbol{i}}$ as
$J_{X}(\boldsymbol{i})=J_{Y}(\boldsymbol{i})=\frac{1}{\sqrt{2}}J_{Z}(\boldsymbol{i})=1+i\sqrt{2\pi G}(\tilde{\alpha}_{\boldsymbol{i}}+\tilde{\alpha}^{\dagger}_{\boldsymbol{i}})-\sqrt{2\pi G}(\tilde{\beta}_{\boldsymbol{i}}+\tilde{\beta}^{\dagger}_{\boldsymbol{i}}).$ (24)
Following a similar procedure as in [18] one can show that the lattice of tunnelling fermions and bosons shown in Fig. 1 (Left) gives rise, in its low energy limit, to a (2+1)(2+1)D Dirac fermion coupled to massive metric fluctuations, as given in (16). While it is straightforward to see the continuum limit of (20) in momentum space leads to Eq. (10), the fermionic Hamiltonian requires some extra analysis. First, one should notice that the Fermi points of the free Hamiltonian ($J_{X}=J_{Y}=J_{Z}/\sqrt{2}=1$) read
$\boldsymbol{P}_{\pm}=\mp\left(\frac{\pi}{2\sqrt{2}},\sqrt{2}\pi\right)$ (25)
Second, expanding the Hamiltonian (22) in momentum space around these Fermi points yields
$H_{\text{f-b}}=\sum_{\boldsymbol{p},\boldsymbol{q}}\Psi^{+\dagger}_{\boldsymbol{p}}\,h^{+}(\boldsymbol{p},\boldsymbol{q})\,\Psi^{+}_{\boldsymbol{q}}+\sum_{\boldsymbol{p},\boldsymbol{q}}\Psi^{-\dagger}_{\boldsymbol{p}}\,h^{-}(\boldsymbol{p},\boldsymbol{q})\,\Psi^{-}_{\boldsymbol{q}},$ (26)
where
$\Psi^{\pm}_{\boldsymbol{p}}=\begin{pmatrix}a^{\pm}_{\boldsymbol{p}}\\ b^{\pm}_{\boldsymbol{p}}\end{pmatrix},\qquad h^{\pm}(\boldsymbol{p},\boldsymbol{q})=\mp A_{\boldsymbol{p}-\boldsymbol{q}}\sigma^{x}q_{x}\mp C_{\boldsymbol{p}-\boldsymbol{q}}\sigma^{x}q_{y}-B_{\boldsymbol{p}-\boldsymbol{q}}\sigma^{y}q_{y}-D_{\boldsymbol{p}-\boldsymbol{q}}\sigma^{y}q_{x},$ (27)
with $\sigma^{x}$ and $\sigma^{y}$ being Pauli matrices, and
$A_{\boldsymbol{k}}=1-\sqrt{2\pi G}(\tilde{\alpha}_{\boldsymbol{k}}+\tilde{\alpha}_{-\boldsymbol{k}}^{\dagger}),\quad B_{\boldsymbol{k}}=1+\sqrt{2\pi G}(\tilde{\alpha}_{\boldsymbol{k}}+\tilde{\alpha}_{-\boldsymbol{k}}^{\dagger}),\quad C_{\boldsymbol{k}}=D_{\boldsymbol{k}}=-\sqrt{2\pi G}(\tilde{\beta}_{\boldsymbol{k}}+\tilde{\beta}_{-\boldsymbol{k}}^{\dagger}).$ (28)
Lastly, back to position space, defining the four-fermion field $\psi$ and the gamma matrices $\gamma^{0}$ and $\gamma^{i}$ as
$\psi=\begin{pmatrix}a^{+}\\ b^{+}\\ b^{-}\\ a^{-}\end{pmatrix},\quad\gamma^{0}=\begin{pmatrix}0&-\boldsymbol{1}\\ \boldsymbol{1}&0\end{pmatrix},\quad\gamma^{i}=\begin{pmatrix}0&\sigma^{i}\\ \sigma^{i}&0\end{pmatrix}$, (29)
allows one to write $H_{\text{f-b}}$ in Dirac form $\int d^{2}x\;\psi^{\dagger}h\,\psi$. When expressing the modes $\tilde{\alpha}$ and $\tilde{\beta}$ in terms of the metric perturbations as given in Eq. (9), the single-particle Hamiltonian $h$ matches precisely (17).
3 Dynamical evolution of a spin-$1/2$ due to geometry fluctuations
In the previous section we presented the lattice Hamiltonian that effectively encodes the dynamics of a Dirac field coupled to fluctuating geometric background. A proposal of how to realise a similar lattice model with optical lattices and a mixture of bosonic and fermionic atoms was given in [18]. Here we want to consider only the evolution of the Dirac fermion’s spin due to the fluctuating background. The minimal subsystem of the lattice that is required to encode and monitor this evolution comprises two fermionic modes that encode the spin of the particle and the two bosonic eigenmodes that couple resonantly to the fermionic ones. Such a minimal system provides an analytically tractable way to predict the time evolution of the system. Moreover, it lends itself for quantum simulations with optical cavity technology as we shall see in this section.
3.1 Spin-$1/2$ coupled to quantum geometry fluctuations
3.1.1 Minimal model
We can isolate the dynamics that the spin-2 field induces on the spin-$1/2$ of the Dirac particle by considering a single fermionic unit cell with $\boldsymbol{i}={\mathbf{0}}$. Omitting the site label, the fermion-boson Hamiltonian reduces to
$H_{\text{f-b}}=J_{Z}a^{\dagger}b+J_{Z}^{\dagger}b^{\dagger}a,$ (30)
where the tunnelling coupling $J_{z}$ is controlled by the bosonic fields in the following way
[…]
We construct a minimal model by considering the Hamiltonian (30) together with the Hamiltonian for gravitational fluctuations in momentum space in the continuum (13), emulating a single spin degree of freedom. This implies that the mode $\tilde{\alpha}=\tilde{\alpha}_{\boldsymbol{i}={\mathbf{0}}}$ is Fourier transformed as
$\tilde{\alpha}=\int\frac{d^{2}k}{2\pi}\tilde{\alpha}(\boldsymbol{k})$ (32)
and similarly for $\tilde{\beta}$. Using relation (11) that express $\tilde{\alpha}$ and $\tilde{\beta}$ in terms of the eigenmodes $\alpha$ and $\beta$, the full Hamiltonian of the system takes the form
$H=\sqrt{2}\bigg(1-\sqrt{\frac{G}{\mu\pi}}\int d^{2}k\big(\beta(\boldsymbol{k})+$ $\beta^{\dagger}(\boldsymbol{k})-i\alpha(\boldsymbol{k})-i\alpha^{\dagger}(\boldsymbol{k})\big)\bigg)a^{\dagger}_{\mathbf{0}}b_{\mathbf{0}}$ (33)
$+\mu\,\int d^{2}k\bigg[\alpha^{\dagger}(\boldsymbol{k})\alpha(\boldsymbol{k})+\beta^{\dagger}(\boldsymbol{k})\beta(\boldsymbol{k})\bigg]+{\rm h.c.}.$
Adopting polar coordinates $k_{x}=k\cos\theta$, and $k_{y}=k\sin\theta$, and expanding the bosonic operators $\alpha$ and $\beta$ in Fourier modes, $\alpha(\boldsymbol{k})=\sum_{n}\alpha_{n}(k)e^{in\theta}$, leads to
$H=\sqrt{2}a^{\dagger}b+\int_{0}^{\infty}kdk\Bigg[2\pi\mu\left(\alpha^{\dagger}_{0}\alpha_{0}+\beta^{\dagger}_{0}\beta_{0}\right)-2\sqrt{\frac{\pi G}{\mu}}\big(\beta_{0}+\beta^{\dagger}_{0}-i\alpha_{0}-i\alpha^{\dagger}_{0}\big)a^{\dagger}b\Bigg]+{\rm h.c.},$ (34)
where we have neglected the Fourier modes $\alpha_{m}(k)$ and $\beta_{m}(k)$ for $m\neq 0$ since they decouple from the fermonic modes. Out of all the possible momenta, the dominant terms in this Hamiltonian are the ones for which the resonant condition $\sqrt{2}\approx 2\pi k\mu$ holds, which defines a circle in momentum space of radius $k_{R}=\frac{1}{\sqrt{2}\pi\mu}$. Dropping the operator labels for the bosonic modes, the Hamiltonian can be rewritten as
$H=\sqrt{2}\sigma^{x}+\sqrt{2}\left(\alpha^{\dagger}\alpha+\beta^{\dagger}\beta\right)+g_{\text{s-b}}\left[\left(\beta+\beta^{\dagger}\right)\sigma^{x}+\left(\alpha+\alpha^{\dagger}\right)\sigma^{y}\right],$ (35)
where $\sigma^{x}=a^{\dagger}b+b^{\dagger}a$, $\sigma^{y}=-i\left(a^{\dagger}b-b^{\dagger}a\right)$ perform the rotation of a spin encoded on the two sites $a$ and $b$ of the unit cell $\boldsymbol{i}$} at half filling, as shown in Fig. 1. Note that the resonance condition makes the self-interaction strengths of the bosonic and fermionic parts to be the same. We also defined the effective spin-boson coupling
$g_{\text{s-b}}=-\frac{\sqrt{2G}}{\sqrt{\pi}\mu^{3/2}}.$ (36)
This minimal system of a spin-$1/2$ particle and two bosonic modes offer an ideal platform to investigate the dynamical evolution of a particle’s spin when coupled to a gravitational field.
3.2 Dynamical evolution of spin
Hamiltonian (35) determines the evolution of a particle’s spin due to its coupling with quantum fluctuations of the gravitational field. Depending on the strength of their coupling this interaction can lead to a variety of time evolutions, including coherent oscillations of spin rotations and fluctuations in the bosonic populations or a fast decoherence of the spin due to the strong gravitational fluctuations.
[…]
where the subscript $f$ denotes the fermionic state, and the subscripts $\alpha$ and $\beta$ denote the two bosonic modes. The time evolution of this state
[…]
for Hamiltonian (35) will in general give oscillatory dynamics between the spin and the bosons. Let us consider the case where $g_{\text{s-b}}=0$: Since the bosonic modes start in the ground state and they are decoupled from the spin, there are no dynamics in the bosonic populations. If the spin is also prepared in an eigenstate of $\sigma^{x}$ then no evolution will take place, while any other spin state will cause it to rotate due to the gravitational background classical field.
Let us now consider the case where $g_{\text{s-b}}\neq 0$. Even if the spin is prepared in the $x$-direction it will evolve as it is coupled non-trivially to the $\beta$-modes. Note that the evolutions corresponding to any other initial spin state will be asymmetrical due to the $\sigma^{x}$ self-energy of the spin. Due to the coupling, the spin-$x$ population is exchanged with the $\beta$-mode population and its oscillation amplitude will decay over time.
To monitor the time-dependent populations of the bosonic modes and fermionic state we define
[…]
where $i=x,y,z$ for the Pauli spin operators.
In the following we will numerically investigate the time evolution for various coupling strengths. Note that $g_{\text{s-b}}$ depends on $G$ and $\mu$. For simplicity we take $\mu=1$, while $G$ is varied such that the regimes $G\ll\mu$, $G\sim\mu$ and $G\gg\mu$ are all considered. In the following section we determine numerically the time-evolution of the spin populations for different coupling regimes and different initial spin states. In the numerical simulation we have to introduce a cut-off for the size of the otherwise infinite dimensional bosonic matrices $\alpha$ and $\beta$. Our numerical analysis shows that increasing the matrix size from $N=14$ to $N=15$ only changes the bosonic populations by at most $0.5\%$. We therefore consider $N=14$ to be a good approximation for these bosonic matrices in our simulation.
3.2.1 Preparation in the spin-$x$ direction
Figure 2: Time evolution of the spin and bosonic populations for weak coupling $G=0.05$ ($\mu=1$), where the spin–boson interaction is much weaker than the bosonic self-interaction. The spin is initially prepared along the $x$-axis. The spin-$x$ population (solid) exhibits coherent exchange with the bosonic modes (dashed/dotted), while spin-$y$ and spin-$z$ remain unpopulated.
Figure 3: Spin-$x$ population dynamics as a function of the coupling $G$ for the system initialized with spin aligned along the $x$-axis ($\mu=1$; see Hamiltonian (35)). At $G=\pi$, the coupling equals $\sqrt{2}$ (the bosonic self-interaction strength). (a) Spin-$x$ population (horizontal axis: time, vertical axis: coupling $G$ on log scale). The white vertical line marks $G=\pi$, roughly separating the coherent from the incoherent time evolutions. The dashed black curve corresponds to the first peak for $G=0.05$ shown in Fig. 3(b). (b) Spin-$x$ dynamics for increasing coupling strength $G$ (spin-$y$ is shown but does not have dynamics). Stronger coupling leads to increased complexity in the spin dynamics.
To study the interplay between spin and bosonic modes, we initialise the system with the spin pointing along the $x$-axis and track its evolution under Hamiltonian (35). In the weak coupling regime ($G\ll 1$), where the spin–boson coupling is much smaller than the bosonic self-interaction strength, the spin-$x$ population exhibits regular, quasi-coherent oscillations, as shown in Fig. 2. These oscillations reflect the reversible exchange of excitation between the spin and the bosonic modes, with spin-$y$ and spin-$z$ remaining unpopulated due to the initial spin alignment and the decoupled dynamics. The frequency of the oscillations is controlled by the coupling $g_{\text{s-b}}$.
The small high-frequency fluctuations, visible on top of the main oscillations, are due to the self-interactions of the bosons and the spin, with coupling $\sqrt{2}$, in (35). As the coupling strength $G$ increases, these micro-oscillations become more pronounced and ultimately dominate the dynamics, disrupting the coherent oscillation pattern. Fig. 3(a) illustrates how the spin-$x$ population evolves as a function of both time and coupling $G$ (plotted on a logarithmic scale). For $G\leq 0.1$, where the self-interaction dominates, the spin-$x$ population exhibits near-perfect revivals, indicating weak backaction from the bosonic environment. At $G=\pi$ (indicated by the white horizontal line in the figure), the spin–boson coupling matches the bosonic self-interaction ($g_{\text{s-b}}=\sqrt{2}$), marking a crossover to a more complex regime. Beyond this point, revivals disappear, and the spin dynamics become increasingly irregular. For strong coupling ($G\geq 10$), the spin rapidly decoheres and fails to return to its initial value.
This transition is more clearly visualised in Fig. 3(b), which shows time slices of the spin-$x$ dynamics for several values of $G$. As $G$ increases, two effects become prominent: (i) the emergence of fast oscillations superimposed on the slower envelope and (ii) a decay in the amplitude of the envelope itself, signalling decoherence due to entanglement between the spin and the bosonic gravitational modes.
These results show that the gravitational bosonic environment acts as an effective decohering bath for the spin, with its strength controlled by the coupling parameter $G$. The absence of full revivals at large $G$ underscores the irreversibility of the dynamics in this regime, which arises from the increased entanglement between the spin and the bosonic degrees of freedom.
3.2.2 Preparation in the spin-$y$ or spin-$z$ direction
The cases where the spin is initially aligned along the $y$- or $z$-axis are similar up to a relabelling of spin components. Consequently, we present results only for the case of initial preparation along the $y$-direction. Fig. 4 displays the time evolution of the system for various coupling strengths $G$, focusing on the spin and bosonic population dynamics.
Fig. 4(a) shows the full population dynamics for weak coupling $G=0.05$. Initially, the spin is fully aligned along the $y$-axis, with zero population in spin-$x$ and spin-$z$. As time evolves, rapid oscillations develop between the spin-$y$ and spin-$z$ populations, reflecting precession in the yyzz plane. The envelope of this oscillation exhibits slower modulations, corresponding to the gradual exchange of population with the bosonic modes. Meanwhile, the spin-$x$ population grows from zero and tracks this envelope, acting as a mediator between the spin subspace and the bosonic environment. The bosonic modes themselves remain nearly identical throughout the evolution due to the symmetry of their coupling.
To highlight how the spin coherence degrades with increasing coupling strength, Fig. 4(b) plots the time evolution of the spin-$x$ and spin-$y$ populations for various values of $G$. At small $G$, clear oscillatory patterns are observed, with near-revivals of the initial spin-$y$ population. As $G$ increases, these revivals fade and are replaced by irregular micro-oscillations and a decay of the overall population envelope, signaling decoherence induced by stronger spin–boson entanglement. The system transitions from regular to chaotic dynamics around $G=\pi$, where the spin–boson coupling becomes comparable to the bosonic self-interaction. Note that the spin-$z$ population, although not shown in Fig. 4(b), closely follows the spin-$y$ dynamics with a relative phase shift of $\pi/2$, consistent with the rotation symmetry of the model in the yyzz plane.
Figure 4: Spin and bosonic population dynamics for a spin initially aligned along the $y$-axis ($\mu=1$; see Hamiltonian (35)). The spin-$z$ population is not shown but closely follows spin-$y$ with a $\pi/2$ phase shift. (a) Population dynamics for $G=0.05$. The spin-$y$ and spin-$z$ populations exhibit rapid oscillations with slowly varying envelopes. The spin-$x$ grows from zero and mediates coupling to the bosonic modes, which remain nearly identical. (b) Time evolution of the spin-$x$ (yellow line) and spin-$y$ (green envelope) populations for different coupling strengths $G$. As $G$ increases, coherence is lost and the system transitions to chaotic dynamics. At $G=\pi$, the spin–boson coupling equals the bosonic self-interaction strength.
3.2.3 Bloch sphere analysis
To characterise the spin’s evolution under its coupling with the bosonic modes, we define the Bloch vector $\boldsymbol{B}(t)=\big(\langle\sigma^{x}(t)\rangle,\;\langle\sigma^{y}(t)\rangle,\;\langle\sigma^{z}(t)\rangle\big)$, where expectation values are taken over the full system state. Figures 2, 3 and 4 display both the trajectories of the Bloch vector and the populations of the bosonic modes for spins initially prepared along different axes.
When the spin is prepared in the xx‑direction, $B_{y}$ and $B_{z}$ remain effectively zero. Hence the spin‐$x$ population (i.e. the probability of measuring “spin up in the xx‑basis”) coincides with the length of the Bloch vector $|\boldsymbol{B}|=\sqrt{\langle\sigma^{x}\rangle^{2}+\langle\sigma^{y}\rangle^{2}+\langle\sigma^{z}\rangle^{2}}$. In contrast, if the spin is initially along $y$, all three components evolve nontrivially. In Figure 5 (especially panel 5(a)) we show the Bloch‐sphere trajectory: the spin vector’s direction undergoes oscillation and its magnitude decays over time. This decay in $|\boldsymbol{B}|$ signals the decoherence of the spin subsystem due to entanglement with the bosonic modes. Under strong coupling (e.g. $G=10$) the decay is rapid. After an initial drop, the magnitude tends to stabilise on average, with residual micro‐oscillations reflecting finite‐size effects of the bosonic sector.
To explore how decoherence scales with the coupling strength, we plot the time‐averaged spin‐vector magnitude in Figure 5(b) as a function of $G$. For small couplings, oscillations of the Bloch vector are regular and revivals of $|\boldsymbol{B}|$ are visible, although the average magnitude is reduced compared to the ideal isolated spin. As $G$ increases, the dynamics becomes more irregular (“chaotic” in appearance), revivals are suppressed, and the time‐averaged magnitude is much smaller. For very large coupling values we even observe a slight increase in the average amplitude compared to immediate large decays, which we attribute to the fact that only a finite number of bosonic modes are resonantly coupled — the system is not truly in the infinite‐bath limit, so complete decoherence (i.e. $|\boldsymbol{B}|\to 0$) does not strictly occur.
A useful comparative reference is the well-known “spin‐boson model”, which describes a two‐level system coupled to an environment of harmonic oscillators. In that model, dephasing and loss of coherence (seen as shrinking Bloch‐vector length) are standard features and have been studied both analytically and numerically. For example, Orth et al. [22] studied nonperturbative decoherence in driven spin‐boson systems, observing suppression of coherence and decay of Bloch‐vector magnitude as coupling increases.
Figure 5: Evolution of the Bloch vector $\boldsymbol{B}(t)=(\langle\sigma^{x}\rangle,\langle\sigma^{y}\rangle,\langle\sigma^{z}\rangle)$ for a spin initially prepared along the $y$-axis ($\mu=1$; dynamics governed by the Hamiltonian in eq. (35)). (a) The spin trajectory on the Bloch sphere for coupling $G=10$. The vector precesses primarily in the yyzz plane while its length gradually decreases, indicating decoherence due to entanglement with the bosonic modes. (b) Time-averaged Bloch vector length as a function of coupling strength $G$. For illustration, the vector length $|\boldsymbol{B}(t)|$ is also shown as a function of time, for two representative couplings, $G=0.05$ and $G=100$ (showing the time-averaged value as a horizontal line), highlighting the transition from weak to strong decoherence.
3.3 Optical cavity quantum simulation
The minimal model Hamiltonian in Eq. (35) describes a spin-$1/2$ particle interacting with two bosonic modes, which is well suited for experimental implementation using cavity quantum electrodynamics (cavity-QED) platforms. In particular, a single atom coupled to a high-finesse optical cavity supporting two orthogonal cavity modes offers a highly controllable realisation of such dynamics that possess long coherence times, suitable for the realisation of the desired time evolutions [23, 24].
[…]
The atom-cavity coupling can be engineered to produce terms proportional to $\sigma^{x}$ and $\sigma^{y}$ via appropriately chosen laser drives and cavity detunings. For example, in the strong-coupling regime, the cavity and atomic transitions hybridise to form an anharmonic ladder of dressed states, as described by the Jaynes–Cummings Hamiltonian [25]. This Hamiltonian naturally includes the effective spin-boson coupling terms of the form $\left(\alpha+\alpha^{\dagger}\right)\sigma^{y}$ and $\left(\beta+\beta^{\dagger}\right)\sigma^{x}$, as required in Eq. (35).
A concrete realisation can be achieved by simultaneously driving the atom and the cavity modes with appropriately tuned lasers. By choosing the laser amplitudes and detunings carefully, one can selectively couple each cavity mode to a specific atomic transition. One cavity mode may be chosen to drive spin rotations around the $x$-axis, while the other mediates rotations around the $y$-axis. The strength of these interactions is directly related to the effective spin-boson coupling parameter $g_{\text{s-b}}$, which in this setup can be tuned through the atom-cavity coupling strength, the detuning, and the drive amplitudes. Notably, such selective coupling strategies have been successfully employed to realise single- and two-photon blockades [25], as well as atom-photon state mapping and deterministic single-photon generation [26].
[…]
4 Conclusions
We have introduced and analysed a minimal model for spin dynamics in a fluctuating quantum background, consisting of a single spin-$1/2$ system coupled to two self-interacting bosonic modes. This toy model captures essential features of matter interacting with a dynamical spacetime environment, while remaining amenable to exact diagonalisation and controlled numerical study.
By preparing the spin in different initial orientations and varying the strength of its coupling to the bosonic modes, we uncovered a rich interplay between coherence, entanglement, and effective decoherence. For weak coupling, the spin exhibits near-periodic oscillations and partial revivals, indicating reversible dynamics with limited backaction from the bosonic modes. As the coupling increases, these revivals are gradually suppressed, giving way to fast, irregular oscillations and decay of the spin population envelope. This is a clear signature of decoherence driven by entanglement with the bosonic gravitational environment.
Despite its simplicity, the model reveals a sharp transition from coherent to chaotic behaviour as the coupling becomes comparable to the bosonic self-interaction strength. This crossover can be interpreted as an effective loss of unitarity from the perspective of the spin, even though the global evolution remains unitary. In this sense, the bosonic sector acts as a gravitational bath, and the spin behaves as a quantum probe of its structure.
We also showed that this model is not purely theoretical: it can be directly realised in current cavity-QED platforms using a single atom coupled to a bimodal optical cavity. In such setups, the spin corresponds to atomic hyperfine states, while the bosonic modes are implemented, e.g. by orthogonally polarised cavity photons with tunable self-interactions. The dynamical regimes discussed in this work, including the onset of decoherence, are therefore accessible with existing experimental tools.
Our findings suggest that minimal quantum simulators can provide valuable insights into effective matter–gravity interactions in a fully quantum regime. Beyond the minimal instance, our full lattice model gives rise gravitational fluctuations of a similar type as those appearing in the emergent gravitational theory that governs the low-energy behaviour of fractional quantum Hall liquids [8, 9], thus providing a controllable platform to explore geometric responses and quantum analogues of curved spacetime in strongly correlated systems. Extensions of this model, such as extending to $(3+1)$ dimensions, adding more bosonic modes, or including measurement backaction, could shed further light on how quantum information behaves in fluctuating or emergent spacetimes.
Acknowledgments
M.S. research projects are supported by the National Research Foundation, Singapore through the National Quantum Office, hosted in A*STAR, under its Centre for Quantum Technologies Funding Initiative (S24Q2d0009). M.S. research is supported by the Ministry of Education, Singapore under the Academic Research Fund Tier 1 (FY2022, A-8000988-00-00). P.S.-R. has been supported by a Young Scientist Training Program (YST) fellowship at the Asia Pacific Center for Theoretical Physics (APCTP) through the Science and Technology Promotion Fund and the Lottery Fund of the Korean Government. P.S.-R. has also been supported by the Korean local governments in Gyeongsangbuk-do Province and Pohang City. This work was partially supported by the EPSRC grants EP/Z533634/1 and UKRI1337:Anyons24.